# 3D Feature Distillation with Object-Centric Priors

### v4

@@ Abstract
Figure 1: Visualization of 3D features (middle), back-projected 2D features (left) and user query similarity heatmaps (right), for previous SOTA point-cloud feature distillation method OpenScene and our DROP-CLIP. OpenScene fuses pixel-wise 2D features across all views with average pooling, leading to grounding failures, segmentation imprecisions and fuzzy object boundaries. Our method resolves these issues by employing object-centric priors to fuse object-level 2D features in 3D instance masks with semantics-informed view selection.
@@ 2 Multi-View Tabletop Objects Dataset
S3DIS [25] indoor - RGB-D,3D $13$ $6$-
Replica [26] indoor - RGB-D,3D $88$- --
@@ 3.2 Employing Object-Centric Priors
Let $\left\{S_{v}^{2D}\in\{0,1\}^{N\times H\times W}\right\}_{v=1}^{\mathcal{V}}$ be view-aligned 2D instance-wise segmentation masks for each scene, where $N$ the total number of scene objects. We aggregate the 2D masks to obtain $S^{3D}\in\{1,\dots,N\}^{M}$, such that for each point $i$ we can retrieve the corresponding object instance $n_{i}=S^{3D}_{i}$.
@@ 3.4 Open-Vocabulary 3D Segmentation
Semantic segmentation In this task, the queries correspond to an open-set of textual prompts $Q=\{\mathbf{q}_{k}\}_{k=1}^{\mathcal{K}}$ describing $\mathcal{K}$ semantic classes. A class for each point $\hat{Y}\in\{1,\dots,\mathcal{K}\}^{M}$ is given by : $\hat{Y}=\texttt{argmax}_{k}\;\texttt{cos}(\hat{Z}^{3D},\mathbf{q}_{k})$.
Referring segmentation Here the user provides an open-vocabulary query $\mathbf{q}^{+}$ referring to a particular object instance, and optionally a set of negative prompts $Q^{-}\in\mathbb{R}^{N^{-}\times C}$, which in practise can be initialized from an open-set as above or with canonical phrases (e.g. ‘object’, ‘thing’ etc.) [13]. Similarity scores are converted to probabilities: $\mathcal{P}=\texttt{softmax}\left(\frac{1}{\gamma}\cdot\texttt{cos}(\hat{Z}^{3D},\,[\textbf{q}^{+},Q^{-}]^{T})\right)$, where $\gamma$ a temperature hyper-parameter and $\mathcal{P}=[\boldsymbol{\rho}^{+},\mathcal{P}^{-}]$ probabilities of positive matching $\boldsymbol{\rho}^{+}\in\mathbb{R}^{M}$ and negative matching $\mathcal{P}^{-}=[\boldsymbol{\rho}^{-}_{1},\dots,\boldsymbol{\rho}^{-}_{N^{-}}]\in\mathbb{R}^{M\times N^{-}}$ respectively. The final 3D segmentation is given by $\hat{S}_{i}=\left(\boldsymbol{\rho}^{+}_{i}>\texttt{max}_{j}\,\mathcal{P}^{-}_{i,j}\right)$ , or by thresholding $\boldsymbol{\rho}^{+}$ with a fixed threshold $s_{thr}$ (see ablations in Appendix A.3)
@@ 4.2 Open-Vocabulary 3D Segmentation Results in MV-TOD
For multi-view, we compare our trained model with OpenScene [12] and OpenMask3D [6] methods, where the full point-cloud from all $73$ views is given as input. We note that for these baselines we obtain the upper-bound 3D features as before, as we observed that our trained model already outperforms them, so we refrained from also distilling features from baselines. For single-view, we feed our network with partial point-cloud from projected RGB-D pair, and compare with 2D baselines MaskCLIP [8] and OpenSeg [9] (see implementation details in Appendix A.4). Our model slightly outperforms the OpenMask3D upper bound baseline in the multi-view setting ($+1.18\%$ in referring and $+2.57\%$ in semantic segmentation), while significantly outperforming 2D baselines in the single-view setting ($>30\%$ in both tasks). Importantly, single-view results closely match the multi-view ones ($\sim-4.0\%$), suggesting that DROP-CLIP indeed learns view-independent features. See Appendix A.5 for more qualitative comparisons with baselines.
@@ 4.3 Generalization to Novel Domains / Tasks
Zero-shot transfer to real-world scenes In this section, we evaluate the zero-shot generalization capability of DROP-CLIP in real-world scenes that contain objects and vocabulary outside the MV-TOD distribution. We test in the validation split of the OCID-VLG [31] dataset, which contains 1249 queries from 165 unique cluttered tabletop scenes. We compare with state-of-the-art 2D CLIP-based baselines LSeg [10], OpenSeg [9] and MaskCLIP [8] for the semantic segmentation task in the single-view setting as before.
@@
Results are presented in Table 4. We find that even though fine-tuned in real data, baselines LSeg and OpenSeg under-perform compared to both MaskCLIP and our DROP-CLIP with a margin of $>10\%$ mIoU, which we attribute to the distribution gap between the fine-tuning dataset ADE20K [40] and OCID scenes. These baselines tend to ground multiple regions in the scene, while MaskCLIP and DROP-CLIP provides tighter segmentations (see Fig. 7). When considering the stricter mAcc75 metric, our approach scores a delta of $7.7\%$ compared to MaskCLIP, suggesting a significant gain in grounding accuracy, especially in cases where the object is heavily occluded. Failures cases were observed in grounding objects that significantly vary in geometry and semantics from the MV-TOD catalog. Please see Appendix A.6.1 for further zero-shot referring segmentation experiments and more qualitative results.
@@
Zero-shot 3D instance segmentation We evaluate the potential of DROP-CLIP for out-of-the-box 3D instance segmentation via clustering the predicted features (see details in Appendix A.6.2). We conduct experiments for both the multi-view setting in MV-TOD, where we compare with Mask3D [41] transferred from the ScanRefer [1] checkpoint provided by the authors, where we feed full point-clouds from 73 views, as well as in OCID-VLG, where we compare with SAM [42] ViT-L model with single-view images. Results are summarized in Table 5. We observe that Mask3D struggles to generalize to tabletop domains, as it has been trained in room layout data with mostly furniture object categories. DROP-CLIP achieves an AP25 of 93.3%, illustrating that the learned 3D features can provide near-perfect instance segmentation in-distribution, even without explicit fine-tuning. When moving out-of-distribution in the single-view setting, we observe that DROP-CLIP achieves mIoU that is competitive with foundation segmentation method SAM ($50.9\%$ vs. $60.1$%). Failure cases include heavily cluttered regions of similar objects with same texture (e.g. food boxes), for which DROP-CLIP assigns very similar features that are identified as a single cluster.
@@ 4.4 Application: Language-guided Robotic Grasping
In this section, we wish to illustrate the applicability of DROP-CLIP in a language-guided robotic grasping scenario. We integrate our method with a 6-DoF grasp detection network [43], which proposes gripper poses for picking a target object segmented by DROP-CLIP. We randomly place 5-12 objects on a tabletop with different levels of clutter, and query the robot to pick a specific object, potentially amongst distractor objects of the same category. The user instruction is open-vocabulary and can involve open object descriptions, attributes, or user-affordances. We conducted 50 trials in Gazebo [44] and 10 with a real robot, and observed grounding accuracy of 84% and 80% respectively, and a final success rate of 64% and 60%. Motion failures were mostly due to grasp proposals for which the motion planning led to collisions. Similar to OCID, grounding failures were due to unseen query concepts and / or instances. Setup and example trials are shown in Fig. 8, more details and visualizations in Appendix A.7 and a video of robot demonstrations is provided as supplementary material.
@@ A.2 Distillation Implementation Details
We use the ViT-L/14@336px variant of CLIP’s vision encoder, which provides features of size $C=768$ from $336\times 448$ image inputs with patch size $14$. We distill with a MinkowskiNet14D [45] sparse 3D-UNet backbone, which consists of 8 sparse ResNet blocks with output sizes of $(32,64,128,256,384,384,384,384)$ and a final $1\times 1$ convolution head to $768$ channels.
@@
To increase the 3D coordinates resolution, we upscale the input point-clouds to ×$\times 10$ and voxelize with original dimension of $d=0.02$ (for feature fusion), and a voxel grid $d=0.05$ for training with the Minkowski framework. To reduce the input dimensionality and speedup training and inference time, we remove the table points via filtering out the table’s 3D mask. 111. During inference, we employ RANSAC to remove table points without access to segmentation masks. We train using AdamW with initial learning rate $3\cdot 10^{-4}$ and cosine annealing to $10^{-4}$ over $300$ epochs, and a weight decay of $10^{-4}$. In each ResNet block, we include sparse batch normalization layers with momentum of $0.1$. We train using two RTX 4090 GPUs, which takes about 4 days. Following  [12, 6] we use spatial augmentations such as elastic distortion, horizontal flipping and small random translations and rotations. We also employ color-based augmentation such as chromatic auto-contrast, random color translation, jitter and hue saturation translation. To better emulate partial views with greater diversity, we train with full point-clouds but further add a per-object blob removal augmentation method that removes consistent blobs of points from each object instance. After training for $300$ epochs, we fine-tune our obtained checkpoint on only partial point-clouds from randomly sampled views for each scene in our dataset. We experimented with several auxiliary losses to reinforce within-object feature similarity, such as supervised contrastive loss [69] as well as KL triplet loss [70], but found that they do not significantly contribute to convergence compared to using only the main cosine distance loss. See Table 7 for a full overview of training and augmentation hyper-parameters.
@@ A.3 Extended Multi-View Feature Fusion Ablations
$\boldsymbol{\rho^{+}}>\mathcal{P}^{-}$ scene 73.7 77.4 73.0 69.8
$\boldsymbol{\rho^{+}}>\mathcal{P}^{-}$ canonical 53.4 57.4 52.6 49.7
$\boldsymbol{\rho^{+}}>\mathcal{P}^{-}$ all 30.8 31.0 31.0 30.8
@@
Inference strategies As discussed in Sec. 3.4 there are two methods for performing referring segmentation inference: (a) selecting all points with higher probability for positive vs. maximum negative prompt $\boldsymbol{\rho}^{+}>\mathcal{P}^{-}$, or (b) thresholding $\boldsymbol{\rho}^{+}$ with a hyper-parameter $s_{thr}$. Additionally, we compare the final referring segmentation performance based on the negative prompts used at test time: (a) prompts from object instances within the scene, (b) prompts from all dataset object instances (similar to semantic segmentation task), (c) fixed canonical phrases {“object”, “thing”, “texture”, “stuff”[13], and (d) no negative prompts (-), where we threshold the raw cosine similarities with the positive query. Results in Table 10. We observe that thresholding provides better results than the first method when the right threshold is chosen, a result which we found holds also for our distilled model. A high threshold of $0.95$ was found optimal for upper bound experiments, while a threshold of $0.7$ for our distilled model, although we further fine-tuned it for zero-shot and robot experiments (see Sec. A.6). Regarding negative prompts, as expected, providing in-scene negatives gives the best results, with a significant delta from canonical ($7.6\%$), all ($7.9\%$) and no negatives ($12.6\%$). However, we observe that even without such prior, the performance is still competitive, even when entirely skipping negative prompts.
@@ A.6.1 Zero-Shot Referring Segmentation Experiments
Table 11: Zero-shot eferring segmentation results in OCID-VLG and REGRAD datasets

### v5

@@ Abstract
Figure 1: Visualization of 3D features (middle), back-projected 2D features (left) and query similarity heatmaps (right), for OpenScene and our DROP-CLIP. OpenScene fuses pixel-wise 2D features with average pooling, leading to grounding failures and fuzzy object boundaries. Our method tackles such issues using object-centric priors to fuse object-level 2D features in 3D instance masks with semantics-informed view selection.
@@ 2 Multi-View Tabletop Objects Dataset
S3DIS [25] indoor - RGB-D,3D $13$ 66 -
Replica [26] indoor - RGB-D,3D $88$ - -
@@ 3.2 Employing Object-Centric Priors
Let $\left\{S_{v}^{2D}\in\{0,1\}^{N\times H\times W}\right\}_{v=1}^{\mathcal{V}}$ be view-aligned 2D instance-wise segmentation masks for each scene, where $N$ the total number of scene objects, provided from the training dataset. We aggregate the 2D masks to obtain $S^{3D}\in\{1,\dots,N\}^{M}$, such that for each point $i$ we can retrieve the corresponding object instance $n_{i}=S^{3D}_{i}$.
@@ 3.4 Open-Vocabulary 3D Segmentation
[…]
Referring segmentation Here the user provides an open-vocabulary query $\mathbf{q}^{+}$ referring to a particular object instance, and optionally a set of negative prompts $Q^{-}\in\mathbb{R}^{N^{-}\times C}$, which in practise can be initialized from an open-set as above or with canonical phrases (e.g. ‘object’, ‘thing’ etc.) [13]. Similarity scores are converted to probabilities: $\mathcal{P}=\texttt{softmax}\left(\frac{1}{\gamma}\cdot\texttt{cos}(\hat{Z}^{3D},\,[\textbf{q}^{+},Q^{-}]^{T})\right)$, where $\gamma$ a temperature hyper-parameter and $\mathcal{P}=[\bm{\rho}^{+},\mathcal{P}^{-}]$ probabilities of positive matching $\bm{\rho}^{+}\in\mathbb{R}^{M}$ and negative matching $\mathcal{P}^{-}=[\bm{\rho}^{-}_{1},\dots,\bm{\rho}^{-}_{N^{-}}]\in\mathbb{R}^{M\times N^{-}}$ respectively. The final 3D segmentation is given by $\hat{S}_{i}=\left(\bm{\rho}^{+}_{i}>\texttt{max}_{j}\,\mathcal{P}^{-}_{i,j}\right)$ , or by thresholding $\bm{\rho}^{+}$ with a fixed threshold $s_{thr}$ (see ablations in Appendix A.3)
@@ 4.2 Open-Vocabulary 3D Segmentation Results in MV-TOD
For multi-view, we compare our trained model with OpenScene [12] and OpenMask3D [6] methods, where the full point-cloud from all $73$ views is given as input. We note that for these baselines we obtain the upper-bound 3D features as before, as we observed that our trained model already outperforms them, so we refrained from also distilling features from baselines. For single-view, we feed our network with partial point-cloud from projected RGB-D pair, and compare with 2D baselines MaskCLIP [8] and OpenSeg [9] (see implementation details in Appendix A.4). Our model slightly outperforms the OpenMask3D upper bound baseline in the multi-view setting ($1.18\%$ in referring and $2.57\%$ in semantic segmentation), while significantly outperforming 2D baselines in the single-view setting ($30\%$ in both tasks). Importantly, single-view results closely match the multi-view ones ($\sim-4.0\%$), suggesting that DROP-CLIP indeed learns view-independent features. See Appendix A.5 for more qualitative comparisons with baselines.
@@ 4.3 Generalization to Novel Domains / Tasks
Zero-shot transfer to real-world scenes In this section, we evaluate the zero-shot generalization capability of DROP-CLIP in real-world scenes that contain objects and vocabulary outside the MV-TOD distribution. We test in the validation split of the OCID-VLG [31] dataset, which contains 1249 queries from 165 unique cluttered tabletop scenes. We compare with 2D CLIP-based baselines LSeg [10], OpenSeg [9] and MaskCLIP [8] and popular 2D grounding method GroundedSAM [40] for the semantic segmentation task in the single-view setting as before.
@@
Results are presented in Table 4. We find that even though fine-tuned in real data, baselines LSeg and OpenSeg under-perform compared to both MaskCLIP and our DROP-CLIP with a margin of $10\%$ mIoU, which we attribute to the distribution gap between the fine-tuning dataset ADE20K [41] and OCID scenes. These baselines tend to ground multiple regions in the scene, while MaskCLIP and DROP-CLIP provides tighter segmentations (see Fig. 7). When considering the stricter mAcc75 metric, our approach scores a delta of $7.7\%$ compared to MaskCLIP, suggesting a significant gain in grounding accuracy, especially in cases where the object is heavily occluded. Failures cases were observed in grounding objects that significantly vary in geometry and semantics from the MV-TOD catalog. Please see Appendix A.6 for further zero-shot experiments, comparisons with modern NeRF/3DGS methods and more qualitative results.
@@
Zero-shot 3D instance segmentation We evaluate the potential of DROP-CLIP for out-of-the-box 3D instance segmentation via clustering the predicted features (see details in Appendix A.6.2). We conduct experiments for both the multi-view setting in MV-TOD, where we compare with Mask3D [42] transferred from the ScanRefer [1] checkpoint provided by the authors, where we feed full point-clouds from 73 views, as well as in OCID-VLG, where we compare with SAM [43] ViT-L model with single-view images. Results are summarized in Table 5. We observe that Mask3D struggles to generalize to tabletop domains, as it has been trained in room layout data with mostly furniture object categories. DROP-CLIP achieves an AP25 of 93.3%, illustrating that the learned 3D features can provide near-perfect instance segmentation in-distribution, even without explicit fine-tuning. When moving out-of-distribution in the single-view setting, we observe that DROP-CLIP achieves mIoU that is competitive with foundation segmentation method SAM ($50.9\%$ vs. 60.160.1%). Failure cases include heavily cluttered regions of similar objects with same texture (e.g. food boxes), for which DROP-CLIP assigns very similar features that are identified as a single cluster.
@@ 4.4 Application: Language-guided Robotic Grasping
In this section, we wish to illustrate the applicability of DROP-CLIP in a language-guided robotic grasping scenario. We integrate our method with a 6-DoF grasp detection network [44], which proposes gripper poses for picking a target object segmented by DROP-CLIP. We randomly place 5-12 objects on a tabletop with different levels of clutter, and query the robot to pick a specific object, potentially amongst distractor objects of the same category. The user instruction is open-vocabulary and can involve open object descriptions, attributes, or user-affordances. We conducted 50 trials in Gazebo [45] and 10 with a real robot, and observed grounding accuracy of 84% and 80% respectively, and a final success rate of 64% and 60%. Motion failures were mostly due to grasp proposals for which the motion planning led to collisions. Similar to OCID, grounding failures were due to unseen query concepts and / or instances. Example trials are shown in Fig. 8, more details in Appendix A.7 and a robot demonstration video is provided as supplementary material.
@@ A.2 Distillation Implementation Details
We use the ViT-L/14@336px variant of CLIP’s vision encoder, which provides features of size $C=768$ from $336\times 448$ image inputs with patch size $14$. We distill with a MinkowskiNet14D [46] sparse 3D-UNet backbone, which consists of 8 sparse ResNet blocks with output sizes of $(32,64,128,256,384,384,384,384)$ and a final $1\times 1$ convolution head to $768$ channels. To increase the 3D coordinates resolution, we upscale the input point-clouds to $\times 10$ and voxelize with original dimension of $d=0.02$ (for feature fusion), and a voxel grid $d=0.05$ for training with the Minkowski framework. To reduce the input dimensionality and speedup training and inference time, we remove the table points via filtering out the table’s 3D mask. 111. During inference, we employ RANSAC to remove table points without access to segmentation masks. We train using AdamW with initial learning rate $3\cdot 10^{-4}$ and cosine annealing to $10^{-4}$ over $300$ epochs, and a weight decay of $10^{-4}$. In each ResNet block, we include sparse batch normalization layers with momentum of $0.1$. We train using two RTX 4090 GPUs, which takes about 4 days. Following  [12, 6] we use spatial augmentations such as elastic distortion, horizontal flipping and small random translations and rotations. We also employ color-based augmentation such as chromatic auto-contrast, random color translation, jitter and hue saturation translation. To better emulate partial views with greater diversity, we train with full point-clouds but further add a per-object blob removal augmentation method that removes consistent blobs of points from each object instance. After training for $300$ epochs, we fine-tune our obtained checkpoint on only partial point-clouds from randomly sampled views for each scene in our dataset. We experimented with several auxiliary losses to reinforce within-object feature similarity, such as supervised contrastive loss [70] as well as KL triplet loss [71], but found that they do not significantly contribute to convergence compared to using only the main cosine distance loss. See Table 7 for a full overview of training and augmentation hyper-parameters.
@@ A.3 Extended Multi-View Feature Fusion Ablations
$\bm{\rho^{+}}>\mathcal{P}^{-}$ scene 73.7 77.4 73.0 69.8
$\bm{\rho^{+}}>\mathcal{P}^{-}$ canonical 53.4 57.4 52.6 49.7
$\bm{\rho^{+}}>\mathcal{P}^{-}$ all 30.8 31.0 31.0 30.8
@@
Inference strategies As discussed in Sec. 3.4 there are two methods for performing referring segmentation inference: (a) selecting all points with higher probability for positive vs. maximum negative prompt $\bm{\rho}^{+}>\mathcal{P}^{-}$, or (b) thresholding $\bm{\rho}^{+}$ with a hyper-parameter $s_{thr}$. Additionally, we compare the final referring segmentation performance based on the negative prompts used at test time: (a) prompts from object instances within the scene, (b) prompts from all dataset object instances (similar to semantic segmentation task), (c) fixed canonical phrases {“object”, “thing”, “texture”, “stuff”[13], and (d) no negative prompts (-), where we threshold the raw cosine similarities with the positive query. Results in Table 10. We observe that thresholding provides better results than the first method when the right threshold is chosen, a result which we found holds also for our distilled model. A high threshold of $0.95$ was found optimal for upper bound experiments, while a threshold of $0.7$ for our distilled model, although we further fine-tuned it for zero-shot and robot experiments (see Sec. A.6). Regarding negative prompts, as expected, providing in-scene negatives gives the best results, with a significant delta from canonical ($7.6\%$), all ($7.9\%$) and no negatives ($12.6\%$). However, we observe that even without such prior, the performance is still competitive, even when entirely skipping negative prompts.
@@ A.6.1 Zero-Shot Referring Segmentation Experiments
Table 11: Zero-shot referring segmentation results in OCID-VLG and REGRAD datasets
@@ A.6.2 Zero-Shot 3D Instance Segmentation Experiments
Integrating spatial object priors via segmentation masks when fusing multi-view features grants separability in the embedding space. To illustrate that, we conducted zero-shot instance segmentation with our DROP-CLIP by directly applying DBSCAN clustering in the output feature space. In our experiments, we use the vanilla implementation of DBSCAN from scikit-learn package and set $\epsilon=0.01,\;\texttt{min\_samples}=2$ for DROP-CLIP and $\epsilon=0.01,\;\texttt{min\_samples}=276$ for OpenScene respectively. We observed that the points that belong to the same object instance are close to each other in the feature space, while significantly differ from the points that belong to other instances. We visualize several examples in Fig. 16, where we also conduct a t-SNE visualization to demonstrate the instance-level separability in the DROP-CLIP feature space.
A.6.3 Comparisons with SfM methods
Method Modality Num. Train Segm. Results
Views Time Model Loc. Sem.Segm.
LERF SfM 171 112.5 min. - 84.8 45.0
LangSplat SfM 171 37.5 min. SAM 88.1 65.1
SemanticGaussians SfM 171 >2 hrs. SAM 89.8 -
LSeg RGB 1 0 - 33.9 21.7
DROP-CLIP RGB-D 1 0 - 66.1 39.1
Table 12: Localization accuracy (%) and 3D semantic segmentation mIoU (%) on the ‘teatime’ scene of LERF dataset. We report number of views, training time and whether / which external models are needed to obtain the representation. Training times are converted to v100 hours from reported numbers in corresponding papers.
@@ A.6.3 Comparisons with SfM methods
We replicate the setup of the localization task from LERF [13] and the semantic segmentation task from LangSplat [24] for the ‘teatime’ scene of the LERF dataset. Results are presented in Table 12, where numbers for representative baselines LSeg [10], LERF [13], LangSplat [24] and Semantic Gaussians [75] are taken from corresponding papers. To signify the aforementioned differences in scope, in our table we also report number of views and training time required to obtain the representation (converted in v100 hours from time reported in corresponding papers) and whether / which external segmentation models (e.g. SAM [43]) is needed during test-time to deal with the ‘patchyness’ issue. The above demonstrate the practical benefits of our approach compared to SfM methods, as mentioned before, working from single-view, real-time performance, zero-shot application and no need for external segmentors. Regarding test results, we find that DROP scores lower to SfM baselines in both task variants, but significantly outperforms LSeg, which is the only other zero-shot baseline. The performance margin between DROP and object-centric 3DGS methods LangSplat and SemanticGaussians is significant, albeit the fact that these methods require SAM at test-time to inject the segmentation priors, whereas DROP doesn’t. This gap is justified when considering that our approach is zero-shot and didn’t have access to the 171 training scenes like the SfM baselines, as well as that the dataset queries are often referring to object parts (e.g. hooves, bear nose etc.), which DROP has not been designed for. Qualitative visualizations of DROP in the LERF scene are given in Fig. 17.
Figure 17: Visualizations of partial point-clouds, 3D DROP-CLIP features (PCA) and similarity heatmaps for three different queries in the ‘teatime’ scene of LERF dataset.